Instructions: Solve each equation by the elimination method.
What is the solution for 4x+2y=3
y=-x

Answer by asha(30)

4X +2Y =3 ---- EQN.1
X =-Y => X+Y =0---- EQN.2
MULTIPLY EQN.2 BY 4
4X + 4Y =0---- eQN.3
-2Y = 3 => Y=- 3/2 (EQN 3 - EQN 1)
PLUGGING THE VALUE OF Y IN EQN 2 WE GET
X= -Y = -(-3/2) =3/2
SOLUTION IS X=3/2 & Y=-3/2
GOOD LUCK!!!